OpenStudy (aravindg):

prove \int\limits^{\pi/2}_{0} \sqrt{1-\sin x)/(1+sinx)dx}=1

OpenStudy (aravindg):

yep

OpenStudy (aravindg):

i think dx is outside

OpenStudy (aravindg):

\[\int\limits\limits^{\pi/2}_{0} \sqrt{1-\sin x)/(1+sinx)}dx=1\]

OpenStudy (apoorvk):

\(LHS=\int\limits^{\pi/2}_{0} \sqrt{(1-\sin x)/(1+sinx)}dx\) \(=\large\int\limits^{\pi/2}_{0} \sqrt{\frac{1 - \sin^2 x}{(1+\sin x)^2}}dx\) \( =\large\int\limits^{\pi/2}_{0} \frac{\cos x}{(1+\sin x)}dx\) Let \((1+\sin x) = t\), then, \(\cos x dx = dt\) So, \( \large\int\limits^{\pi/2}_{0} \frac{dt}{(t)} = \large [\ln t ]^{\pi/2}_{0}= [ln(1+\sin x)]^{\pi/2}_{0}\) = \(\large ln2 - ln1 = ln2\)

OpenStudy (aravindg):

hw is that =1 ?

OpenStudy (anonymous):

@apoorvk , just need to change your limits of integration....

OpenStudy (apoorvk):

@dpaInc but then I switched back to 'x' terms, so why would we convert the limits?

OpenStudy (anonymous):

hmm... ur right.

OpenStudy (apoorvk):

So so so... where are we wrong? Or rather, who's wrong - us or the book? -.-

OpenStudy (anonymous):

the book is wrong... wolfram says it's apoors answer.

OpenStudy (apoorvk):

Aye! :D